1, Efficient Markets Hypothesis (L02,CFA1) you invest $10,000 in the market at the beginning of the year, and by the end of the year your account is worth $15,000. During the year the market return was 10 percent . Does this mean that the market is inefficient?.

2,Geoffrey Henley, a professor of finance, states: The capital market is efficient. I dont know why anyone would bother devoting their time to following individual stock and doing fundamental analysis. the best approach is to buy and hold a well diversified portfolio of stocks. Do you agree? Why or why not?
Answer #1

(Question 1)

Actual return on my investment = [(15,000 - 10,000) / 10,000] x 100% = (5,000/10,000) x 100% = 50%

If market is inefficient, I will be able to earn abnormal returns where

Abnormal return = Actual return on my investment - Market return

In this case,

Abnormal return = 50% - 10% = 40% > 0

Since I could earn abnormal return, it means that the market is inefficient.

NOTE: As per Chegg's Policy, 1st question is answered.